OpenStudy (anonymous):

1) What is the common ratio of the following geometric sequence? 5, 15, 45,…

OpenStudy (kaederfds):

r=3

OpenStudy (anonymous):

What are the next four terms in the following geometric sequence? 5, 15, 45

OpenStudy (anonymous):

I just need help on the last fourth term.

OpenStudy (anonymous):

T4= ar^n-1 = 5*3^4-1 = 5*3^3 = 5*27 T4 = 135 T5= ar^n-1 = 5*3^5-1 = 5*3^4 = 5*81 T5 = 405 T6= ar^n-1 = 5*3^6-1 = 5*3^5 = 5*243 T6= 1215

OpenStudy (kaederfds):

5,15,45 a, ar, ar^2, ar^3,...